OpenStudy (anonymous):

Find f(x) if y = f(x) satisfies \frac(dy)(dx) = 45 yx^(14) and the y -intercept of the curve y = f(x) is 3 .

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= 45yx ^{14}\]

OpenStudy (anonymous):

isn't that stated in the proble,?

OpenStudy (anonymous):

that's the question...

OpenStudy (anonymous):

if it is... then it is separable and you can integrate

OpenStudy (anonymous):

just do implicit differentiation?

OpenStudy (anonymous):

no... separate and integrate...\[\frac{ dy }{ dx }= 45yx^{14} \Rightarrow \frac{ dy }{ y }=45x^{14}dx\]

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

OH yeah now I remember. Thank you!

OpenStudy (anonymous):

the initial condition will help you determine what the constant should be.

OpenStudy (anonymous):

yeah I forgot to separate my variables, that was really helpful!